OpenStudy (anonymous):

Find the least common denominator (LCD) for the fractions. and A. 12 B. 4 C. 16 D. 48

OpenStudy (anonymous):

Which fractions?

OpenStudy (anonymous):

1/12 and 3/4

OpenStudy (anonymous):

plz help ill give medals

OpenStudy (anonymous):

/12

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

can you help with 4 more questions or nah

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Choose the missing number to make equivalent fractions. 1/4 = _/16 A. 2 B. 4 C. 6 D. 8

OpenStudy (anonymous):

It is B.

OpenStudy (anonymous):

Because 4 is a quarter of 16.

OpenStudy (anonymous):

Which number can replace the variable to make these fractions equivalent? n/4=15/20 = A. 2 B. 3 C. 4 D. 5

OpenStudy (anonymous):

It is B.

OpenStudy (anonymous):

Any more?

OpenStudy (anonymous):

Which is a correct comparison of the fractions? 15/18 and 25/30 A. 15/18 < 25/30 B. 15/18 > 25/30 C. 15/18 = 25/30

OpenStudy (anonymous):

help asap